Total probability formula and Bernoulli formula is used for solving
1. There exists 2 student’s group with 30 people in each of them. There are 5 students in the first
group who have the mark “5” on anatomy, 15 students – “good” marks. From the second group
only 3 students have “5” and 10 students – “good”. One student will be chosen from these 2
groups for taking part in the conference. Find out the probability that he will have the mark “5” on
anatomy.
1
Expert's answer
2020-11-04T16:36:45-0500

Common number of students:

n=30*2=60

number of students with mark 5:

N=5+3=8

p=8/60=0.133